OpenStudy (anonymous):

find the equation of the tangent line to y = tan^2(3x) at x= pi/4

OpenStudy (misty1212):

\[y'=2\tan(x)\sec^2(x)\] then you get the slope if you plug in \(\frac{\pi}{4}\)

OpenStudy (anonymous):

oh , so I do the derivative and then plug in the x they give me?

OpenStudy (solomonzelman):

why are you doing the derivative for the user??? also it is wrong!!

OpenStudy (misty1212):

yes it is wrong, i am sorry i did not pay attention to the \(3x\) insides

OpenStudy (solomonzelman):

the slope of this line is: \(\large\color{black}{ f\prime(\pi/4) }\) @misty1212 stop doing the poster work please!

OpenStudy (misty1212):

yes, take the derivative to get a formula for the slope, then plug in the number and you will get the slope

OpenStudy (solomonzelman):

Okay, @ilovereyvis_x3 do you know how to obtain the derivative of this function?

OpenStudy (anonymous):

some what I know that d/dx (tanx) = sex^2x but here tan is ^2 power , so no I do not

OpenStudy (solomonzelman):

I am going to re-write it, and you will tell me if you know how to proceed. Show your wrk though, please. \(\large\color{black}{ y = \tan^2(3x) }\) is same as: \(\large\color{black}{ y = (~\tan(3x)~~)^2 }\)

OpenStudy (anonymous):

is it y prime is (sex(3x) * 3 )^2

OpenStudy (anonymous):

omg I mean sec

OpenStudy (solomonzelman):

you are to: 1) apply the power rule to \(\large\color{black}{ (~\tan(3x)~~)^2 }\) as if \(\large\color{black}{ \tan(3x) }\) is just an x. 2) apply the chain rule to the inner part, \(\large\color{black}{ \tan(3x) }\) 2) apply the chain rule to the inner of the inner part, \(\large\color{black}{ 3x }\)

OpenStudy (anonymous):

no sec ^2 (3x) * 3 )^2

OpenStudy (solomonzelman):

can you tell me what a derivative of tan(x) is?

OpenStudy (anonymous):

sec^2x

OpenStudy (solomonzelman):

yes,

OpenStudy (solomonzelman):

tell me what is the derivative of x^2?

OpenStudy (solomonzelman):

(power rule)

OpenStudy (anonymous):

2x

OpenStudy (solomonzelman):

yes, and the derivative of 3x is 3.

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

now, based on these 3 things we will differentiate your function, okay, lets do it together. \(\large\color{black}{ y=~(~\tan(3x)~~)^2 }\)

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

\(\large\color{black}{ y'=2~(~\tan(3x)~~)^{2-1} \times \sec ^2(3x)\times3}\) \(\large\color{black}{ y'=2~(~\tan(3x)~~)^{1} \times \sec ^2(3x)\times3}\) \(\large\color{black}{ y'=2\tan(3x) \times \sec ^2(3x)\times3}\)

OpenStudy (solomonzelman):

see how I am applying the chain rule twice?

OpenStudy (anonymous):

so you do the power rule thingy and move the 2 to the front and then keep what they gave you which is tan(3x) times sec^2(3x) and then chain rule 3

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

cant you multiply 2 and 3?

OpenStudy (anonymous):

so you get 6tan(3x) * sec^2(3x) ?

OpenStudy (solomonzelman):

so you can notice that I first applied the power rule to the inner argument, then \(\large\color{black}{ \sec ^2(3x)}\) was the chain for \(\large\color{black}{ (~\tan(3x)~~)^{2}}\). and then \(\large\color{black}{ 3}\) is the chain rule for \(\large\color{black}{ 3x}\) that the \(\large\color{black}{ \sec ^2(3x)}\) has.

OpenStudy (solomonzelman):

yes, after we found the derivative we CAN multiply 2 and 3, but i was just shoign where the 3 and 2 are coming from.

OpenStudy (solomonzelman):

\(\large\color{black}{ y'=2\tan(3x) \times \sec ^2(3x)\times3}\) becomes, \(\large\color{black}{ y'=6\tan(3x) \times \sec ^2(3x)}\)

OpenStudy (solomonzelman):

will use a different notation, okay? I'll say \(\large\color{black}{ f'(x)=6\tan(3x) \times \sec ^2(3x)}\)

OpenStudy (anonymous):

yeah that is what I got

OpenStudy (solomonzelman):

cool:)

OpenStudy (solomonzelman):

So, now you need to plug in pi/4 for x, (in the derivatve)

OpenStudy (anonymous):

and now I plug in the x ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do I do sec^2 on my calculator

OpenStudy (solomonzelman):

I will use a calculator for this, because I am lazy. Understanding the concept is important, but not calculating big expressions. See how I am plugging it in? http://www.wolframalpha.com/input/?i=6tan%283%CF%80%2F4%29+sec%5E2%283%CF%80%2F4%29

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it is -12?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

so you know that: \(\large\color{black}{ f~'(\pi/4)=-12}\) \(\large\color{black}{x=\pi/4}\) you need to find the y of the point, knowing that \(\large\color{black}{x=\pi/4}\) plug in \(\large\color{black}{\pi/4}\) for x into f(x).

OpenStudy (anonymous):

ok but what will be the equation of the tangent line

OpenStudy (solomonzelman):

find the y-coordinate of the point

OpenStudy (solomonzelman):

(as I said, plug in \(\large\color{black}{\pi/4}\) into the f(x) )

OpenStudy (anonymous):

y prime is out slope and we have out x but how do I find the y cordinate

OpenStudy (anonymous):

oh the one given

OpenStudy (solomonzelman):

we are given that \(\large\color{black}{x=\pi/4}\), so the y-coordinate is going to be \(\large\color{black}{f(\pi/4)}\) correct?

OpenStudy (anonymous):

i got 1

OpenStudy (solomonzelman):

yes 1 is correct.

OpenStudy (solomonzelman):

okay, a point slope formula says: \(\large\color{black}{y-y_1=m(x-x_1)}\)

OpenStudy (anonymous):

y-1=-12(x-pi / 4 ) ?

OpenStudy (solomonzelman):

yes just like this. now simplify it

OpenStudy (anonymous):

how ???

OpenStudy (solomonzelman):

expand the right side

OpenStudy (anonymous):

don't I just leave it like that?

OpenStudy (solomonzelman):

you want it in y-intercept form probably.

OpenStudy (solomonzelman):

or at least I would think you want it in a form of y=mx+b, but I don't know what your teacher might want you to do.

OpenStudy (anonymous):

it just says write the equation , so I think I don't need to simplify

OpenStudy (anonymous):

oh no , my teacher does not ask for it that way

OpenStudy (solomonzelman):

oh, so then you are done. And even if not, you showed knowledge as far as the calculus part of the problem goes.

OpenStudy (anonymous):

I have another one similar to this one but it sounds more confusing and I do not know what to do

OpenStudy (solomonzelman):

Sure. but make a new question for it please.

OpenStudy (anonymous):

by the way thank you for all the help and yes sure I will to give you another medal

OpenStudy (solomonzelman):

I don't need a medal but I might go offline at any second. THere are those who need me and they can call me at any second. You were a joy to work with, yw;)